Journalist Report – February 28th

Crew 206 – 02/28/2019

Sol 4

Author: Auzou Benjamin, Journalist

"What the hell on Earth is going on?”

On their way to White Rock Canyon, the EVA team stumbled upon a Coke bottle. In the middle of a Martian hiking, Norbert, Jérémy, Aurélien, and Cerise are sent back on Earth.

The planet is endangered, with limited resources, and we continue to pollute and threaten her, even if we are all aware of the danger. Here in the MDRS, we are monitoring our water consumption, our trashes and our energy consumption. For now 4 days we have been following a simple, but healthy and sustainable way of life. We realized that our habits are greedy on resources, on water, on energy. And that it is not so complicated to change and adopt better ones. Even if we want to send humans to Mars, or to the Moon, Earth is and will stay our cradle and our home and we all have the duty to protect her!

Countries have the duty to adopt laws to promote ecology and sustainable development, companies have the duty to use wisely their energies and resources, and at our scale with have the duty to adopt a sustainable way of life. We all have this duty for us, for our children and for all humanity.

In the same time in the Hab, Gaspard and I choose to prepare a surprise olive bread for the crew’s lunch. Unfortunately, that first try was very different from the texture of French baguettes, but everybody loved the taste and it was a huge pleasure to feel a little bit like in France.

After lunch, everybody in the Hab chose to take a nap before the beginning of the human factors experiments. Sadly, both versions of TELEOP experiment failed today: the little rover faced connection issues and the computer of the digital version ran out of battery, with a broken charger. This experiment matters a lot for all of us, especially for Jérémy and the rest of the afternoon was very difficult for us.

The good news of the day came from Cerise’s water monitoring. Yesterday, the results seemed threatening. After an exchange with the research team, the comparison database that she was using was the one dedicated to space, and to Thomas Pesquet’s Aquapad experiment on the ISS. With the correct database, the water quality is not perfect, but safe to drink and not infected.

After the disappointment of the afternoon, we all need a distraction. For that, Norbert brought us some fresh carrots from the Green-Hab for the dinner and we decided to watch an episode of Nat Geo’s Mars TV show.

Greenhab Report – February 28th

Crew 206 Green Hab Report – 28-FEB-2019

GreenHab Officer: Norbert Pouzin

Environmental control: Heating and Cooling w/ ambient air (9 hrs 30 min)

Shade Cloth (40%)

Average temperatures:

Low: 17.8°C

High: 29.5°C

Hours of supplemental light: 3 Hours

Daily water usage for crops: 6.4 gallons

Water in Blue Tank: 270.1 gallons

Crops watered at 0930 hours (2.1 gallons) and 1830 hours (4.3 gallons):

TIME NOTED TEMPERATURE(° C ) / HUMIDITY(%)

1) 09.30 AM 21.4° C / 25 %
2) 11.30 AM 26.3° C / 21 %
3) 15:00 PM 26.7° C / 22 %
4) 18:30 PM 22.3° C / 32 %

Changes to crops: I removed the tomatoes and prepared the soil for the next seeds.
I harvested carrots.
Radishes are germinating.
Red lettuces have some problem growing up and are still not red.

Narrative:

After the morning EVA, I harvested some herbs for our lunch and some carrots for our dinner.

In the afternoon, I prepared the soil for the next tomatoes seeds and removed the old ones.

Finally, I monitor the fan in the Greenhab to avoid high heat in the morning. It will now be turned on from 8 to 16, for 15 min every 15 min. It worked pretty well today!!

The door was opened from 0730 to 1700.

Harvest: 8.1 g of basil, 1.3 g of sage, 270.6 g of carrots.

Support/supplies needed: None.

Operations Report – February 28th

Crew 206 Operations Report 28 – FEB – 2019

SOL: 4

Name of person filing report: Gaspard Thieulin

Non-nominal systems: Generator

Notes on non-nominal systems: Radiator hoses and alternator need replacement on the generator. A contractor has been arranged to make these repairs within a few days.

Generator (hours run): Turned off at 0900 (28th Feb), Turned on at 1730 (28th Feb)

Solar— SOC % – Turned off Gen (28th Feb) 59%, Turned on Gen (28th Feb) 62%

Diesel Reading – 3/4 (on a scale of 0 – 1/4 – 1/2 – 3/4 – 1)

Propane Reading – 80 (± 2)

Ethanol Free Gasoline – 7 gallons

Water (Auxillary tank) – Not in use

Water (Static tank) – around 500 gallons (90 %)

Water differential (Static tank) – 14 gallons (± 2 gallons)

Gallons transferred: Not Applicable

Water in GreenHab – 270.1 gallons

Water (loft) – 12 marked level

Static to Loft Pump used – yes

Water Meter: 01416094 units

Toilet tank emptied: No

Deimos rover used: No, still not functional

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Not Applicable

Beginning charge: Not Applicable

Ending charge: Not Applicable

Currently charging: Not Applicable

Opportunity rover used: Used

Hours: 53.0 hours

Beginning charge: 100%

Ending charge: 61%

Currently charging: Yes

Curiosity rover used: Used

Hours: 77.3 hours

Beginning charge: 100%

Ending charge: 67%

Currently charging: Yes

Spirit rover used: Not Used

Hours: 78.4 hours

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Notes on rovers: Nothing to report.

ATV’s Used: N/A

Reason for use: Not used

Oil Added? N/A

ATV Fuel Used: N/A

# Hours the ATVs were Used today: None

Notes on ATVs: None

Hab Car used and why, where? Not used.

Staff Crew Car used and why, where? Off-Campus

General notes and comments:

Summary of internet: Lost Internet access a lot again during the whole day. Additional information directly transferred to Bernard, MDRS IT.

Summary of suits and radios: Nothing to report.

Summary of Hab operations: Nothing to report.

Summary of Green Hab operations: Nothing to report.

Summary of Science Dome operations: Nothing to report.

Summary of RAM operations: Nothing to report.

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: Nothing to report

Questions, concerns, and requests to Mission Support: Nothing more to report.

EVA Report – February 28th

Crew 206 EVA Report 28-Feb-2019
EVA #4

Author: Norbert Pouzin

Purpose of EVA: The EVA has two purposes:

– the first one is scientific. It is to make sure that the deployment
of the LOAC and the weather station is fine. Since the snow has melted, we have to re-calibrate the tripod of the weather station. Also, the batteries of the LOAC have to be changed.

– the second purpose of the EVA is the exploration for the other half of the crew who didn’t perform it today. We will go to White Rock Canyon and explore this area. We will park the rovers on the road 1101, at the closest point from the White Rock Canyon, and then walk to the canyon. As the distance between the MDRS and White Rock Canyon is a little bit smaller than the one we did today going to The Moon, the batteries of the rover might not be a problem (34% battery left for Opportunity, 65% for Curiosity).

However, if it appears that the batteries aren’t strong enough to bring us there, we will explore Kissing Camel Ridge West, parking the rovers on Cow Dung Road 0110, near Robert’s Rock Garden. Like today, we prefer being four for this EVA, so that two of us can work on the tripod and the others on the LOAC and we can be two on
each rover, which is safer than being alone.

Start time: 9:00 am

End time: 11:26 am

Narrative: Today, we went pretty fast on our EVA preparation, so we left the hab at 0900, instead of 0915, as it was planned. After the usual checks (Water tank for Team Red, composed of Aurélien and Jeremy, and rovers states for Team Blue, composed of Cerise and me), we went to the localization of the LOAC and the weather station.

We arrived there at 0915. While Team Red changed the batteries of th LOAC, Team Blue measured the inclination of the weather station, as the melting snow might have moved the tripod. Indeed, the angle had moved a little, so we tried to fix it by turning the screw of the tripod. We left the area of the LOAC at 0931, heading into White Rock Canyon with our rovers, taking care of the batteries, like yesterday. It worked pretty well, since we parked at 0952 on the road near the Canyon, with 90% batteries for Curiosity, and 85% for Opportunity!
We started the exploration, following our scout Jeremy. We went inside the canyon, and then hiked to go above it.

The soil wasn’t as muddy as it was yesterday, so we could easily walk. At 1058 our walk was over, and we came back with our rovers. We arrived at the station at 1120, with more than 60% of batteries on both rovers.

Destination: LOAC and Weather Station location, Road 1101, White Rock Canyon

Coordinates: 518500/425700 (LOAC and Weather Station), 520250/4247600 (White Rock Canyon)

EVA Participants: Norbert Pouzin (GHO, EVA leader), Aurélien Mure (XO), Jeremy Auclair (CMD), Cerise Cuny (HSO).

Roads and routes per MDRS Map: Entrance Road, Cow Dung Road

Mode of travel: Driving and Walking

Sol Summary – February 28th

Crew 206 February 28th, 2019
Sol 4

Summary Title: French bread in the Station

Author: Jeremy Auclair

Mission Status: All nominal

Sol Activity Summary: Muscle aches from the morning workout have finally passed, it feels pretty good. Our sleep patterns are getting better. This morning’s EVA’s purpose was to change the LOAC battery (which went perfectly) and to explore White Rock Canyon. The rovers performed better than yesterday, we managed to arrive at White Rock Canyon with more than 85% of charge. The exploration lasted for about 1h30, the view was exceptional and we took great pictures. The mud is still bothering, but nothing that we can’t overcome. We arrived at the Hab a little tired. For lunch, we ate pasta and we cooked our first bread!

This afternoon I worked on TELEOP again, but I encountered new issues: the computer charger is broken, and we haven’t found an equivalent one. I managed to connect the physical rover to my computer by there is still connection issues. To sum up: both TELEOP versions are currently not working. It was a hard day for me, but the rest of the crew had more luck. Cerise got a response from the BioMérieux lab: our water is not infected (we feared there were some bacteria), it’s safe to drink, but it’s not as clean as tap water. We are still experiencing internet issues, it seems to be coming from the router (I couldn’t connect before 07:08 pm).

Look ahead plan: Tomorrow will be a regular day, we will try to fix some issues on our experiments. The crew is impatient for Sol 6; they will try their space suit for the first time in real conditions.

Anomalies in work: None

Weather: Sunny in the morning, partly cloudy in the afternoon

Crew physical status: Fine

EVA: White Rock Canyon

Reports to be filed: GreenHab, Journalist, Operation, EVA Report, EVA Request

Support Requested: New water filters for the jug

Journalist Report – February 27th

Crew 206 – 02/27/2019

Sol 3

Author : Auzou Benjamin, Journalist

"Martian Excursion"

We are born to explore; since the dawn of humanity, humans never stopped
to explore new places. The first humans settled to all five continents,
the navigators of the XVth century sailed all seas, the astronauts of
the XXth century flew to space. Today was an exploration day on Mars !
The team of the EVA (Aurélien, Norbert, Gaspard and I) rode the rovers
to the Moons area. We walked about an hour on the muddy and slippery
slopes of the White and Grey Moons, following the steps of Norbert, our
scout of the day, under the wise supervision of the EVA leader Aurélien.

To avoid fatigue after lunch, I visited Norbert in the Green-Hab and
helped him taking care of radishes and carrots while some others were
taking a nap. Back in the Hab, Aurélien was waiting for us all to show
us his first night observation : Orion’s Nebula.

This afternoon we tested our own analog spacesuit in the MDRS for the
first time. This project was born in September 2017, when we first heard
about the MDRS and decided to apply to a rotation. We worked since that
moment on the conception and fabrication of this spacesuit. That’s why
we are happy to see this project concretise here. The prototype is far
from what we expected but we are focused on our goal to take some of us
in EVA with this suit and we will make all changes and adjustments
needed to reach it.

Today we had the first results of our strict water monitoring. And this
is a huge improvement compared to the monitoring of crew 189, the last
mission of ISAE-SUPAERO students. For Sol 2, the total consumption is
77.4 liters. The Green-Hab is the first source of consumption with 41%,
then come drinking water with 25%, the flushes with 13%, cooking water
and dishes water with 5% and 6%. The last 8% are unmeasured and come
from leaks or flushes volume uncertainties. In comparison, the average
unmeasured water for crew 189 was 37% of total consumption.

For more than a year, we worked together on the preparation of this
mission. We passed through happy and bad moments, funny talks and long
crisis meetings. But now we are here, on Mars, and we are like a family,
forged by our common passion for space exploration, our interventions
with young students and this mission. What we are achieving now is
exciting and fulfilling !

Sol Summary – February 27th

Crew 206 February 27th 2019
Sol 3

Summary Title : Taking care of the Hab

Author : Jeremy Auclair

Mission Status : All nominal

Sol Activity Summary : Once again we got out of bed at 07:00 am to
workout, but everybody had slept well, we went to bed early yesterday.
Our Dreem Headbands showed that we had an effective time of sleep
ranging from 6h20 and 8h20, I slept more than everyone else ! After
breakfast we prepared the EVA. It was the first EVA without a veteran,
Aurélien was EVA leader. It went very well, Opportunity’s battery
decreased quickly so they went for plan B. They explored the moon areas.
During the EVA Cerise and I cleaned the Hab’s lower and upper deck. I
then took care of the water consumption data : on Sol 2 (yesterday) we
used 77,4L (20,5 Gallons), with 41% for the GreenHab, 25% for drinking,
13% for flushes, 6% for dishes, 5% for cooking, 2% for hygiene (no
showers yet) and 8% of unmeasured consumption. Unmeasured consumption
comes from inaccuracies and small mistakes in counting, this part of
daily consumption should decrease during the mission. After lunch,
everyone worked on their experiments (except me, I took a good nap).
Aurélien, Benjamin, Gaspard and Norbert did some pre-EVA tests on their
space suit, all the vital functions work, but it still needs to be
improved the ergonomic standpoint. Cerise worked on Aquapad to measure
the microbial quality of our water. First results suggest that the water
drinkability is problematic, but it is surely a manipulation or data
interpretation, we have sent a mail to the BioMérieux laboratories to
have an explanation. Before the COMMS, as planned, we participated to the
Virtual Reality experiment.

Look ahead plan : Tomorrow is once again going to be a busy day, we have
4 experiments planned. Tomorrow’s EVA will allow us to train the last
members of the crew.

Anomalies in work : None

Weather : Mostly cloudy

Crew physical status : Fine

EVA : Beige moon area

Reports to be filed : GreenHab, Journalist, Operation, EVA Report, EVA
Request, Astronomy

Support Requested : None

Operations Report – February 27th

Crew 206 Operations Report 27 – FEB – 2019

SOL: 3

Name of person filing report: Gaspard Thieulin

Non-nominal systems: Generator

Notes on non-nominal systems: Radiator hoses and alternator need
replacement on generator. A contractor has been arranged to make these
repairs within a few days.

Generator (hours run): Turned off at 0800 (27th Feb), Turned on at 1840
(27th Feb)

Solar— SOC % – Turned off Gen (24th Feb) 64%, Turned on Gen (24th Feb) 59%

Diesel Reading – 3/4 (on a scale of 0 – 1/4 – 1/2 – 3/4 – 1)

Propane Reading – 21 (± 2)

Ethanol Free Gasoline – 7 gallons

Water (Auxillary tank) – Not in use

Water (Static tank) – around 500 gallons (90 %)

Water differential (Static tank) – 12 gallons (± 2 gallons)

Auxillary to Static tank transfer – No

Gallons transferred: Not Applicable

Water in GreenHab – 276.5 gallons

Water (loft) – 12 marked level

Static to Loft Pump used – yes

Water Meter: 01415960 units

Toilet tank emptied: Yes

Deimos rover used: No, still not functional

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Not Applicable

Beginning charge: Not Applicable

Ending charge: Not Applicable

Currently charging: Not Applicable

Opportunity rover used: Used

Hours: 52.4 hours

Beginning charge: 100%

Ending charge: 34%

Currently charging: Yes

Curiosity rover used: Used

Hours: 76.7 hours

Beginning charge: 100%

Ending charge: 61%

Currently charging: Yes

Spirit rover used: Not Used

Hours: 78.3 hours

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Notes on rovers: Opportunity’s battery decreased faster
under 60%, and lasted way less than Curiosity’s. (for more
detailed battery readings during the EVA, cf EVA report)

ATV’s Used: N/A

Reason for use: Not used

Oil Added? N/A

ATV Fuel Used: N/A

# Hours the ATVs were Used today: None

Notes on ATVs: None

Hab Car used and why, where? Used for Staff

Staff Crew Car used and why, where? Off Campus

General notes and comments:

Summary of internet: Lost Internet access a lot again this
morning. Did a software reset on the modem. Working
without interruption now.

Summary of suits and radios: Suit #9 Fixed, currently charging. Suit #2
experienced fogging. Checked it, seems ok

Summary of Hab operations: Hotfixed the lamp bulb in upper
deck. It is stable now. Inspection revealed the plastic bulb is broken on
several loactions.

Summary of Green Hab operations: Nothing to report

Summary of Science Dome operations: Nothing to report

Summary of RAM operations:

Nothing to report.

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing more to report.